OpenStudy (nerdy_3000):

Can someone please provide me with a quick answer to this: The naturalist said that 50% of the 2100 plants in the park are flowering. Find the number of plants that are flowering?

OpenStudy (anonymous):

1050 plants is the answer 50/100=2100/x solver for x=1050

OpenStudy (anonymous):

1050